A yo-yo factory has fixed operating costs of $575,000 per year. In addition to the fixed costs, the cost to produce one yo-yo is $1.30. The yo-yo company sells the yo-yos to a distributor for $5.75 each. How many yo-yos must be sold in a year to earn a profit of at least $140,000?
Enter a whole number of yo-yos.
At least__________yo-yos

Answer by addingup(3677)

575,000 + 1.30x + 140,000 = 5.75x
1.30x = -715,000 + 5.75x
-4.45x = -715,000 = 160,675
-----------------------------
Check:
160,675 * 5.75 = 923,881.25 total revenue. Deduct expenses:
923,881.25 - (575,000 + 1.30(160,675)) = 140,000 Correct